Graph the function y=tan(1/4)x in the interval from -2pi to 2pi.
Answer by lwsshak3(11628)

Graph the function y=tan(1/4)x in the interval from -2pi to 2pi.
..
Standard form for tangent function: y=tan(Bx-C), with period=π/B, phase-shift=C/B
For given function:
y=tan(x/4-0)
B=1/4
period=π/(1/4)=4π
1/4 period=π
Graphing the function
on the x-axis make tick marks at -2π, -π, 0, π, 2π (tick marks spaced 1/4 period apart)
x=-2π
y=tan(Bx)=tan(-2π/4)=tan(-π/2)=undefined (asymptote)
x=-π
y=tan(-π/4)=-1
x=0
y=tan(0)=0
x=π
y=tan(π/4)=1
x=2π
y=tan(2π/4)=tan(π/2)=undefined (asymptote)
You now have the following points to graph the given tan function:
(-2π, asymptote), (-π,-1), (0,0), (π,1), (2π, asymptote)